A Complete Bibliography of Publications in Computational Statistics & Data Analysis (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

$17.75 [85]. $28.50 [95]. $29.95/£ [52, 16]. 2^{n} [150]. $32.00 [53]. $70.00 [14]. ^{(}R) [83]. th [197]. ^{T M} [75, 80]. ^{X} [12]. A [214]. C_{p} [88]. E [164]. L_{1} [32]. λ [225]. r × c [149].

-optimal [214, 164]. -type [88].

0-444-00658-3 [14]. 0-444-86453-9 [15]. 0-471-10466-3 [52, 16].

11 [90]. 13th [195]. 1989 [197].

28 [252].

6th [170]. accuracy [126, 183, 249]. additive [192]. AGREE [93]. agreement [93]. air [35]. Alessandra [252]. algorithm [157, 223, 214, 40, 59]. algorithms [246, 126, 183, 149]. alone [235]. among [242, 178]. Amsterdam [15]. Analele [243]. Analysis [252, 113, 21, 9, 45, 111, 70, 186, 184, 170, 199, 52, 20, 117, 218, 101, 144, 255, 16, 114, 65, 158, 33, 27, 163, 18, 85]. Analyzing [151]. Announcement [29, 54, 132, 17, 197]. Announcements [147, 153, 159]. Announces [171, 45, 71, 175, 174, 237, 231, 239]. Announcing [241]. annual [195]. ANOVA [249, 216, 221]. ANOVA-designs [221]. APL [27]. Apollo [49]. Apple [69]. application [246, 60]. Applications [134, 85]. applied [26, 148]. approach [145, 165, 89, 144, 41, 216, 18]. approaches [86]. approximants [112]. Approximately [180]. Approximating [97]. approximation [60, 228]. approximations [32]. arrays [220]. Assessing [249]. Assessment [42]. associated [133]. asymmetric [163]. August [186]. Author [55, 120, 160, 205, 250]. AUTOBJ [24]. AUTOBJ/AUTOBOX [24]. AUTOBOX [24]. available [36, 44, 79, 80, 91]. award [103].

B [63, 143, 141, 146, 213]. B-spline [213]. background [192]. Bailey [138, 140]. balanced [137, 138]. bandwidth [224]. based [189, 30, 224, 40, 163]. baseline [151]. basic [136]. basics [85]. Bayes [164]. Bayesian [189, 165, 188, 60]. Behavioral [58]. Behrens [156]. Benchmark [4]. Benchmarking [183]. Bernoulli [96, 136]. best [221]. between [20]. binary [177, 248]. binomial [151]. bivariate [180, 228]. block [137, 215, 138, 164]. BMDP [13, 92, 83, 203, 235, 236, 237]. BMDP82 [36]. Board [56, 57, 162, 207]. Book [53, 52, 15, 14, 85, 16]. booth [239]. Boston [95]. Bounded [222]. box [233]. Brown [63, 143, 141, 146]. Bucuresti [243]. bundling [233]. business [106, 62].

C [85]. C. [186, 63, 143, 141, 98]. Calculating [182]. calculations [249]. calibrated [180]. Call [185, 197]. capabilities [23]. career [21]. Carlo [33, 191]. case [218]. categorical [122]. categories [189]. cell [189]. censored [33]. census [70]. changes [133]. Charlottesville [196]. Chemometrics [16]. Chichester [52, 16]. Chishire [53]. choice [88]. choosing [40]. CICS [109]. Classification [119, 196, 189, 195, 98]. clustering [86]. coefficient [194, 154]. collinearity [179]. collinearity-influential [179]. Combinations [209]. Comment [63, 143, 141]. Comments [224]. communications [118]. comparative [253, 14]. Comparison [86, 20, 135, 168, 33, 219]. component [20, 101]. components [252, 165, 211]. comprehensive [173]. computation [168, 98, 191]. Computational [252]. computations [246]. computer [173, 254, 15, 227]. Computerized [22]. computers [36, 44, 91, 172]. Computing [247, 85, 23, 100, 87, 93, 164, 134]. Computrition [94]. Conference [196, 102, 68, 186, 170]. Confidence [136, 96]. connected [150]. Connectivity [242]. considerations [225]. consulting [71]. contemporary [95]. contingency [63, 141, 125, 2, 144, 149]. continuous [32]. conversational [44]. convex [191]. correlation [113, 124]. covariance [20, 113]. covariate [33]. CP [69]. CP/M [69]. Cray [171]. criteria [209]. critique [138]. Cromp [139]. cross [189]. cross-classification [189]. CT [53]. cycles [62].

dasy [111]. Data [252, 21, 9, 111, 74, 186, 242, 238, 199, 151, 112, 52, 188, 122, 85, 255, 16, 180, 224, 33, 27, 192, 40, 4, 163, 97]. data-based [40, 163]. database [242]. databases [232]. DATATRIEVE [70]. David [15, 85]. DEC [79, 234]. decomposition [187]. definite [182]. degree [227]. densities [60]. density [32, 40]. dependent [165, 248]. design [217, 89, 218, 210]. designing [168]. designs [252, 137, 215, 138, 211, 212, 213, 214, 209, 164, 221, 219]. detection [86]. determination [135]. development [133]. developments [186]. Dfl [15]. Diagnosing [179]. diagnostics [35]. diagrams [22, 166]. dichotomy [225]. dictionary [74]. differential [241]. Din [243]. directions [164]. directories [69]. Discrete [32, 60, 247, 97]. discrimination [42, 145, 144]. disease [133]. dispersion [220]. Display [107, 53]. distributed [227]. distribution [194, 154, 97, 167]. distributions [245, 123, 247, 41, 166, 149]. dose [192]. dose-response [192]. during [246]. Duxbury [85, 95].

econometric [19]. editor [72, 229, 169]. Editorial [1, 56, 57, 162, 207]. Edward [53]. effects [33, 220]. efficiency [221]. Efficient [246, 212]. electronic [50]. elements [182]. Elsevier [14, 67, 128]. enhancements [108]. entries [125]. entropy [30, 247]. entropy-based [30]. equations [76, 241]. equireplicate [212]. Erratum [28, 230]. estimated [166]. estimates [31]. estimating [189, 155, 40]. Estimation [31, 123, 252, 63, 141, 181, 2, 99, 223, 188, 211, 60, 41, 248]. European [185]. eutrophication [218]. Evaluating [227]. evaluation [4]. event [21]. event-history [21]. exact [154, 225, 149, 219]. expanded [172]. expectation [227]. Experimental [252, 211, 217, 89, 213]. experiments [150, 168]. exploratory [85]. Exploring [62]. exponential [223]. EXPORT [74]. extension [247].

F [139, 85]. facilities [66]. facility [74]. factorial [150]. families [223]. fast [157]. feasible [164]. Federation [196]. Fifth [102]. files [74]. Findler [139]. First [197, 21, 45]. fisher [42, 156]. fit [30, 97]. Fitting [192, 61, 59, 157]. fixed [135, 149]. flexible [193, 4]. forecasting [43, 77]. Formula [227]. FORSYS [77]. FORTRAN [193]. four [226]. four-measure [226]. Francis [14]. free [98]. FSCALC [109]. Fuchs [63, 143, 141]. Fully [210]. Fully-sequential [210]. function [191, 155]. functions [32, 227, 61]. Further [141].

gamma [123]. GCOS [80]. general [238, 87]. generalised [61]. Generalized [176, 190, 192, 167]. generators [245]. Genstat [6, 234]. German [21]. GEV [190]. Giovagnoli [252]. given [247]. GLIM [197, 151]. global [59]. Goodman [225]. goodness [30, 97]. goodness-of-fit [30]. graph [233, 108, 175]. graph-in-the-box [233]. Graphics [53, 27, 232]. graphing [166]. group [71]. guide [172]. Gumbel [190].

H [95]. having [189]. heuristic [89, 248]. Hewlett [78]. higher [227]. higher-degree [227]. history [21]. Hoaglin [85]. Holland [15, 14]. homogeneity [5]. Honeywell [80]. honor [186]. Hugh [15]. hypothesis [87].

IAS [116]. IAS-SYSTEM [116]. IBM [92, 73, 69, 129, 127]. IFAC [102]. IFAC/IFORS [102]. IFORS [102]. Illustrating [35]. illustration [21]. image [170]. immunoassay [155]. Implementation [176]. IMPORT [74]. IMPORT/EXPORT [74]. improved [59]. improvement [115]. Inc. [45, 46, 71, 104, 239, 232, 238, 240]. incomplete [137, 138]. increases [233]. index [55, 120, 160, 121, 161, 205, 250, 206, 251]. industrial [52, 255, 16]. inference [186, 180]. influence [226]. influential [179]. information [45, 145, 144, 53]. Informix [231, 232]. instruments [90]. integrated [41]. interactive [117, 43, 27]. interface [70]. interfaces [231, 232]. International [186, 196, 197]. interpolation [217]. interval [136]. intervals [96]. Introduces [198, 46, 104, 235, 232]. introduction [77, 95, 116]. inverse [182]. IRIS [184]. ISBN [52, 15, 14, 16]. isotonic [222]. ISSUE [68]. Italy [197]. iterative [157, 223]. IV [9]. Ivor [14].

J [52]. John [52, 16]. joint [197]. July [197]. June [196].

Keegel [145]. Kermit [118]. known [194, 154]. Koopmans [95]. Kruskal [225]. Kullback [145].

lake [218]. Lambert [95]. latent [176]. latest [203, 237]. launches [67]. least [193]. Letter [229, 169]. Letters [72]. Lewi [52]. library [66, 23, 152]. likelihood [63, 141, 2, 194, 190, 122, 163]. line [22]. LInear [25, 42, 181, 244, 157, 148, 122, 88, 193, 87, 192, 61]. link [61]. LISREL [45]. LOg [25, 157, 122, 88]. log-linear [157, 122, 88]. logistic [42, 167]. lognormal [123, 59]. LOLITA [25]. loosely [137, 138].

M [69]. M. [63, 143, 141, 146]. MA [85]. Macintosh [202]. mainframe [118]. management [9, 106]. Manager [107]. manipulation [227]. manual [254, 15]. margins [149]. market [82]. Mass [26]. Massachusetts [95]. matrices [182, 20, 227]. Matrix [98]. MAVIS [77]. Maximum [122, 63, 141, 2, 247]. maximum-entropy [247]. McGill [115]. mean [252, 194, 211, 154, 191]. means [135, 180]. measure [226]. measurements [133]. measures [178]. median [222]. medians [62]. meeting [185, 195, 197]. meeting-psychometric [185]. member [41]. Menu [75]. merging [212]. method [155, 164]. methods [189, 122, 100, 87, 33]. metric [112]. Micro [118]. microcomputer [83]. microcomputers [126, 34]. MicroMath [175, 174]. microprocessor [26]. Microstat [84]. MicroVax [234]. minicomputers [105]. Minimal [150]. minimax [215]. minimum [145, 144]. Minitab [8]. MINQUE [98]. MINSQ [174]. missing [180]. Mixtures [244]. mobility [21]. Model [189, 35, 190, 88, 248, 133]. Model-based [189]. modelling [25, 197, 112]. models [173, 252, 42, 176, 181, 157, 211, 122, 98, 19, 192, 61, 59, 219]. modified [223, 166].

Modistat [38]. Modulad [204]. Monte [33, 191]. Monte-Carlo [33]. mortality [35]. mosaics [134]. mr [82]. MSUSTAT [117]. multifactor [5]. multimode [176]. multinomial [181]. multiple [168]. Multivariate [112, 255, 184, 190, 60, 87, 114, 133, 52, 16].

N. [139]. NAG [66]. NCGA [239]. NCR [238]. NCR/TOWER [238]. nearest [86]. neighbor [86]. nested [98]. Neuchâtel [186]. nominal [93]. non [235, 215, 148, 212, 192]. non-equireplicate [212]. non-linear [148]. non-randomized [215]. non-statisticians [235]. non-zero [192]. nonlinear [210]. Nonparametric [177]. nonzero [151]. normal [194, 154, 135, 180, 228]. normally [227]. normed [178]. North [15, 14, 85]. North-Holland [15, 14]. note [137, 145, 138]. Number [115, 219]. Numerical [126, 183, 59]. nutritional [94].

observations [179, 18]. obtain [96].

obtained [212]. occur [99]. one [135, 114]. one-way [114]. Optimal [213, 32, 214, 164]. optimality [215, 209]. optimization [59]. optimum [89, 209]. option [233]. Order [167, 221]. ordered [189]. ordinary [241]. orthogonal [220]. OSIRIS [9]. other [86]. outliers [99]. outlying [18]. overview [240, 24, 43]. Oxford [15]. OXYCALC [50].

P [10, 49, 82, 130, 129]. P-STAT

[10, 49, 82, 130, 129]. P-STAT/mr [82]. PACK [24]. Package [17, 121, 161, 206, 251, 115, 48, 184, 254, 47, 15, 117, 93]. packages [34]. Packard [78]. paired [219]. pairs [219]. Paola [252]. papers [185, 197]. Parameter [181, 96, 123, 136, 41, 248, 59]. parametric [61]. path [22]. Paul [52, 85]. PC [73, 92, 73, 69, 118, 129, 127, 236, 198, 233, 199]. PC/AT [129]. PC/XT [129]. Pearson [41]. percentile [31]. personal [172]. planning [106]. plot [226]. point [40]. Poisson [245]. pollution [35]. ported [49, 129]. positive [182]. possible [187]. power [87, 225, 163]. powerful [241]. pp [53, 52, 15, 14, 85, 16, 95]. Practical [87, 96, 216]. practice [255, 52, 16]. prediction [225]. Preface [208]. Premium [99]. Press [53, 85, 95]. prime [36, 91]. principal [20, 101]. printer [22]. Prism [7]. prize [103]. probability [247]. problem [136, 156]. problems [210]. procedure [96, 45, 99, 19, 248]. procedures [210]. product [48]. products [242]. program [50, 76, 184, 173, 193]. programs [67, 237, 176, 254, 15]. project [106]. properties [148]. proportional [157, 223, 221]. proportions [189, 178]. protection [99]. Pro^{T M} [46]. PROTRAN [51]. psychometric [185].

QR [187]. quadratic [42, 219]. quantitative [53]. Quasi [188, 163]. Quasi-Bayesian [188]. quasi-likelihood [163].

R [186, 53, 98]. R. [139, 138, 140]. Random [245, 217, 115]. Randomization [216]. randomized [215]. ranking [47]. RANKSEL [47]. Rao [186, 98]. ratio [194, 190]. Reduction [219, 100]. regression [173, 222, 35, 213, 193, 177, 214, 248, 18]. regressions [244, 187]. regular [223]. Rejoinder [64, 142, 145, 138, 143]. related [245]. relations [20]. relative [221]. release [79, 203, 236, 237, 234]. reliability [165]. reliable [155]. remarks [194]. repeated [133]. repetitive [246]. REPLAY [109]. REPLAY-CICS [109]. replication [177]. Reply [140]. report [121, 161, 206, 251]. report/Review [206, 121, 161, 251]. reports [70]. Research [16, 82, 119]. Response [139, 146, 140, 99, 192]. results [124]. Review [206, 53, 84, 52, 15, 14, 85, 16, 121, 161, 251, 253, 5, 14]. Revue [204]. role [152]. rule [18]. RUNNER [43]. running [62].

SADP [48]. sales [77]. Sample [135, 168, 180, 228]. sampling [218]. SAS [11, 106, 109, 108, 105, 172]. SAS/GRAPH [108]. SAS/OR [109]. SCA [37]. scale [93]. scaling [223]. Scandinavian [170]. scatter [166]. scientific [128]. scientists [199]. Scituate [85]. SCSS^{T M} [44]. Sebastiani [252]. second [221, 196]. selected [182]. selection [47, 19, 41]. selector [224]. September [16]. sequential [210]. Series [16, 44]. sets [188, 4]. several [33]. SGCSA [103]. SHAZAM [173]. ships [236]. SIBYL [43]. SIBYL-RUNNER [43]. Simple [228, 156]. simulation [100]. six [5]. size [135, 168, 228]. small [180]. Smart [104]. SmartForecasts [104]. smoothers [148]. Societies [196]. society [185, 195]. Software [104, 65, 158, 9, 67, 69, 128, 131, 241, 203, 235, 236, 237, 238, 240, 253, 3, 14, 58, 249, 4]. Solo [235]. solution [235]. solutions [82]. solving [241]. Some [194, 20, 148, 124]. Sons [52, 16]. sorting [246]. source [254, 15]. space [86]. space-time [86]. sparse [63, 141, 2]. splicing [148]. spline [213]. spread [5]. spreadsheet [50]. SPSS [45, 46, 12, 74, 79, 71, 73, 90, 80, 91, 78, 118, 131, 171, 231, 232, 198, 233]. SPSS-11 [90]. SPSS-X [74, 79, 80, 91, 78, 171, 131]. SPSS-X^{T M} [45]. SPSS/PC [73, 118, 198, 233]. SPSS/Pro^{T M} [46]. squares [193]. STAN [39]. stand [235]. stand-alone [235]. standard [176]. STAT [10, 49, 51, 130, 129, 240]. STAT/mr [82]. STAT/PROTRAN [51]. STATCAT [15, 83, 254]. STATGRAPHICS [27]. statistic [88, 114]. Statistical [253, 133, 9, 37, 66, 186, 203, 235, 236, 237, 242, 197, 240, 3, 246, 117, 126, 23, 136, 65, 158, 27, 183, 249, 4, 34, 14]. statisticians [235, 152]. Statistics [252, 26, 190, 227, 95, 149, 167]. Statpro [110, 127]. Stein [18]. Stein-rule [18]. Stigler [188]. structural [76]. structures [113]. Studies [16, 100, 58]. study [218]. Subjective [101]. subroutine [152]. subset [187]. support [191]. survey [82, 3, 65, 158, 149, 34]. survival [33]. Switzerland [186]. Sybase [231, 232]. Sygraph [201, 239]. symmetric [182]. SYSTAT [239, 238, 200, 202, 240]. System [105, 26, 9, 44, 45, 37, 12, 94, 111, 172, 165, 43, 27, 24, 116]. systems [79, 77].

TAbles [25, 189, 63, 141, 125, 2, 99, 144, 149]. Targets [199]. technical [70]. Test [220, 194, 190, 154, 156, 225, 149]. testing [87]. tests [5, 30, 216, 97]. Texas [90]. theory [217]. Three [69, 123, 59]. three-parameter [123, 59]. time [193, 100, 86]. time-varying [193]. tool [106, 119, 242]. TOWER [238]. transfers [242]. transformation [163]. transport [74]. treatment [33]. treatments [212]. Trento [197]. TSP [81]. TUFTE [53]. two [237, 99, 165]. two-way [99]. type [88].

unbalanced [98]. Unicos [171]. Universitatii [243]. University [115]. Update [100, 45]. Upgraded [127]. USA [196]. use [163]. user [254, 15, 58]. using [151, 88, 62, 187, 166].

V [139]. V3.0 [198]. VA [196]. value [233]. values [180]. variability [178]. variable [176, 19, 248]. variance [252, 31, 211, 114, 155]. variate [245]. variation [194, 154]. varying [193]. Vax [234, 79, 131]. vector [124]. Velleman [85]. Version [81, 173, 127, 174, 239, 200, 202]. via [193]. visual [53]. VMS [234]. Vol [16]. volume [55, 120, 160, 121, 161, 205, 250, 206, 251].

VS [44].

Wang [44]. wars [240]. way [99, 114]. Weibull [123]. West [21]. width [40]. Wiley [52, 16]. window [40]. within [192]. without [177]. working [21]. workshop [197].

X [131, 74, 79, 80, 91, 78, 171]. XT

[73, 110, 129]. X^{T M} [45].

years [21]. York [15, 14].

Zero [125, 192].

References

Azen:1983:E

[1] Stanley P. Azen. Editorial. Computational Statistics & Data Analysis, 1(??):1–2, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/0167947383900580.

Brown:1983:MLE

[2] Morton B. Brown and Camil Fuchs. On maximum likelihood estimation in sparse contingency tables. Computational Statistics & Data Analysis, 1(??):3–15, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/0167947383900592. See

comments [63, 141].

Francis:1983:SSS

[3] Ivor Francis. A survey of statistical software. Computational Statistics & Data Analysis, 1(??):

17–27, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900609.

Wilson:1983:BDS

[4] S. R. Wilson. Benchmark data sets for the flexible evaluation of statistical software. Computational Statistics & Data Analysis, 1(??): 29–39, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900610.

Games:1983:RSM

[5] Paul A. Games and Glenn S. Wolfgang. A review of six multifactor tests for homogeneity of spread. Computational Statistics & Data Analysis, 1(??):41–52, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900622.

Anonymous:1983:G

[6] Anonymous. Genstat. Computational Statistics & Data Analysis, 1 (??):53, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900634.

Anonymous:1983:P

[7] Anonymous. Prism. Computational Statistics & Data Analysis, 1(??): 54–55, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900646.

Anonymous:1983:Ma

[8] Anonymous. Minitab. Computational Statistics & Data Analysis, 1 (??):55–57, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900658.

Anonymous:1983:OIS

[9] Anonymous. The OSIRIS IV statistical analysis and data management software system. Computational Statistics & Data Analysis, 1 (??):57–60, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738390066X.

Anonymous:1983:PS

[10] Anonymous. P-STAT. Computational Statistics & Data Analysis, 1 (??):60–61, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900671.

Anonymous:1983:Sa

[11] Anonymous. SAS. Computational Statistics & Data Analysis, 1(??): 61–62, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900683.

Anonymous:1983:STS

[12] Anonymous. The SPSS^{X} ^{T M} system. Computational Statistics & Data Analysis, 1(??):63, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900695.

Anonymous:1983:B

[13] Anonymous.  BMDP.  Computa-

tional Statistics & Data Analysis, 1 (??):63–65, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900701.

Friedman:1983:BRB

[14] Herman P. Friedman. Book review: Statistical software: a comparative review: by Ivor Francis, Elsevier North-Holland, New York, 1981 (550 pp., $70.00, ISBN 0-444-00658-3). Computational Statistics & Data Analysis, 1(??):67–68, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900713.

Frane:1983:BRB

[15] James W. Frane. Book review: The computer package STATCAT source programs and user manual: by Hugh David, North-Holland, Amsterdam, New York, Oxford, 1982 (780 pp., Dfl. 240.00, ISBN 0-444-86453-9). Computational Statistics & Data Analysis, 1(??):69–71, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900725.

Lewi:1983:BRB

[16] Paul J. Lewi. Book review: Multivariate data analysis in industrial practice: Chemometrics Research Studies Series Vol. 3 John Wiley & Sons, Chichester, September 1982 (256 pp, $29.95/£13.75, ISBN 0-471-10466-3). Computational Statistics & Data Analysis, 1(??):72, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900737.

Anonymous:1983:PA

[17] Anonymous. Package announcement. Computational Statistics & Data Analysis, 1(??):73, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900749.

Ziemer:1983:SRA

[18] Rod F. Ziemer and Jean-Paul Chavas. A Stein-rule approach to outlying observations in regression analysis. Computational Statistics & Data Analysis, 1(??):75–83, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900750.

Onishi:1983:VSP

[19] Haruo Onishi. A variable selection procedure for econometric models. Computational Statistics & Data Analysis, 1(??):85–95, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900762.

Flury:1983:SRB

[20] Bernhard Flury. Some relations between the comparison of covariance matrices and principal component analysis. Computational Statistics & Data Analysis, 1(??):97–109, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900774.

Andress:1983:FYW

[21] Hans-Jürgen Andress. The first 10 years of a working career: an illustration of event-history analysis with West German mobility data. Computational Statistics & Data Analysis, 1 (??):111–135, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900786.

Huba:1983:CPD

[22] G. J. Huba and A. L. Palisoc. Computerized path diagrams on a line printer. Computational Statistics & Data Analysis, 1(??):137–140, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900798.

Malone:1983:LNC

[23] James M. Malone II. A library of new capabilities in statistical computing. Computational Statistics & Data Analysis, 1(??):141–143, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900804.

Chupalio:1983:OPS

[24] Frances M. Chupalio and David P. Reilly. An overview of the PACK SYS-TEM and AUTOBJ/AUTOBOX. Computational Statistics & Data Analysis, 1(??):145–146, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900816.

Anonymous:1983:LLL

[25] Anonymous. LOLITA (LOg LInear modelling in TAbles). Computational Statistics & Data Analysis, 1 (??):147–149, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900828.

Anonymous:1983:MMA

[26] Anonymous. Mass microprocessor applied statistics system. Computational Statistics & Data Analysis, 1 (??):149–151, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738390083X.

Polhemus:1983:SID

[27] Grace P. Polhemus. STATGRAPHICS: an interactive data analysis and statistical graphics system in APL. Computational Statistics & Data Analysis, 1(??):151–152, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900841.

Anonymous:1983:E

[28] Anonymous. Erratum. Computational Statistics & Data Analysis, 1 (??):153, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900853.

Anonymous:1983:Aa

[29] Anonymous. Announcement. Computational Statistics & Data Analysis, 1(??):155–156, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900865.

Gokhale:1983:EBG

[30] D. V. Gokhale. On entropy-based goodness-of-fit tests. Computational Statistics & Data Analysis, 1(??): 157–165, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900877.

Brown:1983:EVP

[31] Morton B. Brown and Robert A. Wolfe. Estimation of the variance of percentile estimates. Computational Statistics & Data Analysis, 1 (??):167–174, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900889.

Katz:1983:DAC

[32] Darryl Katz. Discrete approximations to continuous density functions that are L_{1} optimal. Computational Statistics & Data Analysis, 1 (??):175–181, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900890.

Nessim:1983:MCC

[33] S. Nessim and D. L. Preston. A Monte-Carlo comparison of several methods for the analysis of censored survival data with treatment and covariate effects. Computational Statistics & Data Analysis, 1(??):183–190, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900907.

Woodward:1983:SSP

[34] Wayne A. Woodward and Alan C. Elliott. A survey of statistical packages on microcomputers. Computational Statistics & Data Analysis, 1 (??):191–200, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900919.

Gibbons:1983:IRD

[35] Diane I. Gibbons and Gary C. Mc-Donald. Illustrating regression diagnostics with an air pollution and mortality model. Computational Statistics & Data Analysis, 1(??):201– 220, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900920.

Anonymous:1983:BAP

[36] Anonymous. BMDP82 available for prime computers. Computational Statistics & Data Analysis, 1(??): 221, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900932.

Anonymous:1983:SSS

[37] Anonymous. The SCA statistical system. Computational Statistics & Data Analysis, 1(??):221– 223, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900944.

Anonymous:1983:Mb

[38] Anonymous. Modistat. Computational Statistics & Data Analysis, 1 (??):224–226, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900956.

Anonymous:1983:Sb

[39] Anonymous.  STAN.  Computa-

tional Statistics & Data Analysis, 1 (??):226–227, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900968.

Sheather:1983:DBA

[40] Simon Sheather. A data-based algorithm for choosing the window width when estimating the density at a point. Computational Statistics & Data Analysis, 1(??):229–238, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738390097X.

Parrish:1983:IAM

[41] Rudolph S. Parrish. On an integrated approach to member selection and parameter estimation for Pearson distributions. Computational Statistics & Data Analysis, 1(??): 239–255, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900981.

Bayne:1983:AFL

[42] C. K. Bayne, J. J. Beauchamp, V. E. Kane, and G. P. McCabe. Assessment of fisher and logistic linear and quadratic discrimination models. Computational Statistics & Data Analysis, 1(??):257–273, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383900993.

Mahmoud:1983:OSR

[43] Essam H. Mahmoud. An overview of the SIBYL-RUNNER interactive forecasting system. Computational Statistics & Data Analysis, 1(??): 275–277, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901007.

Anonymous:1983:SCS

[44] Anonymous. SCSS^{T M} conversational system now available for Wang VS series computers. Computational Statistics & Data Analysis, 1(??): 277–278, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901019.

Anonymous:1983:SIA

[45] Anonymous. SPSS Inc. announces new LISREL procedure in first update of SPSS-X^{T M} information analysis system. Computational Statistics & Data Analysis, 1(??):278–279, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901020.

Anonymous:1983:SII

[46] Anonymous. SPSS Inc. introduces SPSS/Pro^{T M}. Computational Statistics & Data Analysis, 1(??):279, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947383901032.

Edwards:1983:RRS

[47] Howard P. Edwards. RANKSEL: a ranking and selection package. Computational Statistics & Data Analysis, 1(??):279–281, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383901044.

Anonymous:1983:SPP

[48] Anonymous. SADP product package. Computational Statistics & Data Analysis, 1(??):281–282, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383901056.

Anonymous:1983:PSP

[49] Anonymous. P-STAT ported to Apollo. Computational Statistics & Data Analysis, 1(??):282–283, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383901068.

Anonymous:1983:OES

[50] Anonymous. OXYCALC electronic spreadsheet program. Computational Statistics & Data Analysis, 1(??): 283–285, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738390107X.

Anonymous:1983:SP

[51] Anonymous. STAT/PROTRAN. Computational Statistics & Data Analysis, 1(??):285–288, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383901081.

Clark:1983:BRB

[52] Virginia Clark. Book review: Multivariate data analysis in industrial practice: by Paul J. Lewi, John Wiley and Sons, Chichester, 1982 (256 pp., $29.95/ £13.75, ISBN 0-471-10466-3). Computational Statistics & Data Analysis, 1(??):289–290, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901093.

Azen:1983:BRB

[53] Stanley P. Azen. Book review: The visual display of quantitative information: by Edward R. TUFTE Graphics Press, Chishire, CT, 1983 (190 pp., $32.00). Computational Statistics & Data Analysis, 1(??):290–291, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738390110X.

Anonymous:1983:Ab

[54] Anonymous. Announcement. Computational Statistics & Data Analysis, 1(??):292–293, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901111.

Anonymous:1983:AIV

[55] Anonymous. Author index volume 1 (1983). Computational Statistics & Data Analysis, 1(??):294– 295, March 1983. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947383901123.

Anonymous:1983:EB

[56] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 1(??):IFC, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947383900579.

Anonymous:1984:EB

[57] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 2(1):ii, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900276.

Molenaar:1984:BSS

[58] Ivo W. Molenaar. Behavioral studies of the software user. Computational Statistics & Data Analysis, 2 (1):1–12, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900288.

Wingo:1984:FTP

[59] Dallas R. Wingo. Fitting threeparameter lognormal models by numerical global optimization — an improved algorithm. Computational Statistics & Data Analysis, 2(1): 13–25, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738490029X.

Katz:1984:DAM

[60] Darryl Katz and David Z. D’Argenio. Discrete approximation of multivariate densities with application to Bayesian estimation. Computational Statistics & Data Analysis, 2(1): 27–36, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900306.

Scallan:1984:FPL

[61] A. Scallan, R. Gilchrist, and M. Green. Fitting parametric link functions in generalised linear models. Computational Statistics & Data Analysis, 2 (1):37–49, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900318.

Polasek:1984:EBC

[62] Wolfgang Polasek. Exploring business cycles using running medians. Computational Statistics & Data Analysis, 2(1):51–70, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738490032X.

Aston:1984:CMB

[63] Christopher E. Aston and Susan R. Wilson. Comment on M. B. Brown and C. Fuchs, “On maximum likelihood estimation in sparse contingency tables”. Computational Statistics & Data Analysis, 2(1):71–77, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900331. See

Brown:1984:R

[64] Morton B. Brown and Camil Fuchs. Rejoinder. Computational Statistics & Data Analysis, 2(1):79–80, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0167947384900343. See [142].

Neffendorf:1984:SSS

[65] Hugh Neffendorf, David Cable, and Beverly Rowe. Software for statistical and survey analysis. Computational Statistics & Data Analysis, 2 (1):81–94, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900355.

Anonymous:1984:NLS

[66] Anonymous. NAG library statistical facilities. Computational Statistics & Data Analysis, 2(1):95, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900367.

Anonymous:1984:ELS

[67] Anonymous. Elsevier launches software programs. Computational Statistics & Data Analysis, 2(1):95–97, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0167947384900379.

Anonymous:1984:IC

[68] Anonymous. ISSUE conference. Computational Statistics & Data Analysis, 2(1):97, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900380.

Anonymous:1984:TDA

[69] Anonymous. Three directories — for Apple, IBM PC, and CP/M software. Computational Statistics & Data Analysis, 2(1):98–99, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900392.

Anonymous:1984:NTR

[70] Anonymous. New technical reports on census analysis, DATATRIEVE interface. Computational Statistics & Data Analysis, 2(1):99, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900409.

Anonymous:1984:SIA

[71] Anonymous. SPSS Inc. announces new consulting group. Computational Statistics & Data Analysis, 2 (1):100, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900410.

Anonymous:1984:LE

[72] Anonymous. Letters to the editor. Computational Statistics & Data Analysis, 2(1):101–102, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900422.

Anonymous:1984:SPI

[73] Anonymous. SPSS/PC for the IBM PC XT. Computational Statistics & Data Analysis, 2(1):103, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900434.

Anonymous:1984:IEF

[74] Anonymous. IMPORT/EXPORT facility in SPSS-X to transport data and dictionary files. Computational Statistics & Data Analysis, 2 (1):104, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900446.

Anonymous:1984:M

[75] Anonymous. Menu ^{T M}. Computational Statistics & Data Analysis, 2(1):104–105, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900458.

Anonymous:1984:NSE

[76] Anonymous. New structural equations program. Computational Statistics & Data Analysis, 2(1):105– 106, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738490046X.

Mahmoud:1984:ISF

[77] Essam Mahmoud. An introduction to sales forecasting systems: FORSYS and MAVIS. Computational Statistics & Data Analysis, 2 (1):107–108, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900471.

Anonymous:1984:SXH

[78] Anonymous. SPSS-X on Hewlett– Packard. Computational Statistics & Data Analysis, 2(1):108, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/0167947384900483.

Anonymous:1984:NRS

[79] Anonymous. New release of SPSS-X now available on DEC VAX systems. Computational Statistics & Data Analysis, 2(1):108–110, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900495.

Anonymous:1984:SXAa

[80] Anonymous. SPSS-X^{T M} available on Honeywell GCOS. Computational Statistics & Data Analysis, 2 (1):110, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900501.

Anonymous:1984:TV

[81] Anonymous. TSP version 4.0. Computational Statistics & Data Analysis, 2(1):110–112, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900513.

Anonymous:1984:PSM

[82] Anonymous. P-STAT/mr — solutions for market and survey research. Computational Statistics & Data Analysis, 2(1):113, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900525.

Anonymous:1984:BSM

[83] Anonymous. The BMDP StatCat^{(}R) microcomputer. Computational Statistics & Data Analysis, 2(1):113– 116, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900537.

Bechtel:1984:RM

[84] Roy Bechtel. Review of microstat 4.0. Computational Statistics & Data Analysis, 2(1):117–121, June 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900549.

Hill:1984:BRB

[85] Mary Ann Hill. Book review: Applications, basics, and computing of exploratory data analysis by Paul F. Velleman and David C. Hoaglin Duxbury Press, North Scituate, MA (354 pp., $17.75). Computational Statistics & Data Analysis, 2(1):

123–124, June 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900550.

McAuliffe:1984:CNN

[86] Timothy L. McAuliffe and A. A. Afifi. Comparison of a nearest neighbor and other approaches to the detection of space-time clustering. Computational Statistics & Data Analysis, 2(2):125–142, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738490001X.

Muller:1984:PMC

[87] Keith E. Muller and Bercedis L. Peterson. Practical methods for computing power in testing the multivariate general linear hypothesis. Computational Statistics & Data Analysis, 2(2):143–158, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900021.

Jolayemi:1984:CLL

[88] E. Tejumola Jolayemi and Morton B. Brown. The choice of a log-linear model using a C_{p}-type statistic. Computational Statistics & Data Analysis, 2 (2):159–165, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900033.

Findler:1984:HAO

[89] Nicholas V. Findler and Robert F. Cromp. A heuristic approach to optimum experimental design. Computational Statistics & Data Analysis, 2 (2):167–178, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900045.

Anonymous:1984:STI

[90] Anonymous. SPSS-11 for the Texas instruments 990. Computational Statistics & Data Analysis, 2(2): 179, August 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900057.

Anonymous:1984:SXAb

[91] Anonymous. SPSS-X available on prime computers. Computational Statistics & Data Analysis, 2(2):179– 180, August 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900069.

Anonymous:1984:BIP

[92] Anonymous. BMDP on the IBM PC. Computational Statistics & Data Analysis, 2(2):180–182, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900070.

Popping:1984:APC

[93] Roel Popping. AGREE, a package for computing nominal scale agreement.

Computational Statistics & Data Analysis, 2(2):182–185, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900082.

Anonymous:1984:CNS

[94] Anonymous. Computrition nutritional system. Computational Statistics & Data Analysis, 2(2):186– 187, August 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900094.

Nessim:1984:ICS

[95] S. Nessim. An introduction to contemporary statistics: Lambert h. Koopmans duxbury Press, Boston, Massachusetts, 1981 (599 pp., $28.50). Computational Statistics & Data Analysis, 2(2):189, August 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900100.

Alvarez:1984:PPO

[96] Olga Alvarez, Andrzej Matuszewski, and David Sotres. A practical procedure to obtain confidence intervals for the Bernoulli parameter. Computational Statistics & Data Analysis, 2 (3):191–206, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900112.

Zelterman:1984:ADG

[97] Daniel Zelterman. Approximating the distribution of goodness of fit tests for discrete data. Computational Statistics & Data Analysis, 2 (3):207–214, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900124.

Kleffe:1984:MFC

[98] J. Kleffe and B. Seifert. Matrix free computation of C. R. Rao’s MINQUE for unbalanced nested classification models. Computational Statistics & Data Analysis, 2(3):215– 228, November 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900136.

Draper:1984:PPR

[99] N. R. Draper, Irwin Guttman, and J. A. John. Premium and protection of a response estimation procedure for twoway tables when outliers occur. Computational Statistics & Data Analysis, 2 (3):229–236, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900148.

Matloff:1984:UMR

[100] Norman S. Matloff. Update methods for reduction of computing time in simulation studies. Computational Statistics & Data Analysis, 2(3):237– 242, November 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738490015X.

Korhonen:1984:SPC

[101] Pekka J. Korhonen. Subjective principal component analysis. Computational Statistics & Data Analysis, 2 (3):243–255, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900161.

Anonymous:1984:FII

[102] Anonymous. Fifth IFAC/IFORS conference. Computational Statistics & Data Analysis, 2(3):257, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947384900173.

Anonymous:1984:SPA

[103] Anonymous. SGCSA prize award. Computational Statistics & Data Analysis, 2 (3):257–258, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900185.

Anonymous:1984:SSI

[104] Anonymous. Smart Software, Inc. introduces SmartForecasts. Computational Statistics & Data Analysis, 2 (3):259–260, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900197.

Anonymous:1984:SSN

[105] Anonymous. The SAS system now for minicomputers. Computational Statistics & Data Analysis, 2(3):260– 261, November 1984. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900203.

Anonymous:1984:SBP

[106] Anonymous. SAS business planning and project management tool. Computational Statistics & Data Analysis, 2(3):262, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900215.

Anonymous:1984:DM

[107] Anonymous. Display manager. Computational Statistics & Data Analysis, 2 (3):262–263, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947384900227.

Anonymous:1984:SGE

[108] Anonymous. SAS/GRAPH enhancements. Computational Statistics & Data Analysis, 2(3):263–264, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0167947384900239.

Anonymous:1984:SFR

[109] Anonymous. SAS/OR, FSCALC, REPLAY-CICS. Computational Statistics & Data Analysis, 2(3):264, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947384900240.

Anonymous:1984:SX

[110] Anonymous. Statpro XT. Computational Statistics & Data Analysis, 2(3):265, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947384900252.

Anonymous:1984:DDA

[111] Anonymous. dasy — data analysis system. Computational Statistics & Data Analysis, 2(3):266, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947384900264.

Brannigan:1985:MDM

[112] M. Brannigan. Multivariate data modelling by metric approximants. Computational Statistics & Data Analysis, 2 (4):267–278, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900015.

Lee:1985:ACC

[113] Sik-Yum Lee. Analysis of covariance and correlation structures. Computational Statistics & Data Analysis, 2 (4):279–295, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900027.

Nath:1985:NSO

[114] Ravinder Nath and Robert Pavur. A new statistic in the one-way multivariate analysis of variance. Computational Statistics & Data Analysis, 2 (4):297–315, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900039.

Pierchala:1985:IMU

[115] Carl E. Pierchala. An improvement for the McGill University random number package. Computational Statistics & Data Analysis, 2(4):317– 322, February 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900040.

Sonnberger:1985:IIS

[116] Harald Sonnberger. An introduction to the IAS-SYSTEM. Computational Statistics & Data Analysis, 2 (4):323–330, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900052.

Isitt:1985:MIS

[117] John Isitt. MSUSTAT — An interactive statistical analysis package. Computational Statistics & Data Analysis, 2 (4):331–336, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900064.

Anonymous:1985:MMC

[118] Anonymous. Micro to mainframe communications with SPSS/PC and Kermit. Computational Statistics & Data Analysis, 2(4):337, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900076.

Anonymous:1985:CTR

[119] Anonymous. Classification as a tool of research. Computational Statistics & Data Analysis, 2(4):337– 338, February 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900088.

Anonymous:1985:AIVa

[120] Anonymous. Author index volume 2 (1984). Computational Statistics & Data Analysis, 2(4):339– 340, February 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590009X.

Anonymous:1985:PRRa

[121] Anonymous. Package report/review index volume 2 (1984). Computational Statistics & Data Analysis, 2 (4):341, February 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900106.

Haber:1985:MLM

[122] Michael Haber. Maximum likelihood methods for linear and log-linear models in categorical data. Computational Statistics & Data Analysis, 3 (??):1–10, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900532.

Kappenman:1985:ETP

[123] Russell F. Kappenman. Estimation for the three-parameter Weibull, lognormal, and gamma distributions. Computational Statistics & Data Analysis, 3(??):11–23, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900544.

Robert:1985:SRV

[124] P. Robert, R. Cléroux, and N. Ranger. Some results on vector correlation. Computational Statistics & Data Analysis, 3(??):25–32, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900556.

Baker:1985:ZEC

[125] R. J. Baker, M. R. B. Clarke, and P. W. Lane. Zero entries in contingency tables. Computational Statistics & Data Analysis, 3(??):33–45, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900568.

Lesage:1985:NAS

[126] James P. Lesage and Stephen D. Simon. Numerical accuracy of statistical algorithms for microcomputers. Computational Statistics & Data Analysis, 3(??):47–57, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590057X.

Anonymous:1985:UVS

[127] Anonymous. Upgraded version of STATPRO for the IBM PC. Computational Statistics & Data Analysis, 3(??):59, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900581.

Anonymous:1985:ESS

[128] Anonymous. Elsevier scientific software.

Computational Statistics & Data Analysis, 3(??):60–67, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900593.

Anonymous:1985:PSP

[129] Anonymous. P-STAT ported to IBM PC/XT and PC/AT. Computational Statistics & Data Analysis, 3(??):67, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590060X.

Anonymous:1985:PS

[130] Anonymous. P-STAT 8. Computational Statistics & Data Analysis, 3 (??):68–71, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900611.

Anonymous:1985:SXS

[131] Anonymous. SPSS-x software for VAX. Computational Statistics & Data Analysis, 3(??):72–73, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900623.

Anonymous:1985:Aa

[132] Anonymous. Announcement. Computational Statistics & Data Analysis, 3(??):75–76, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900635.

Tango:1985:SMC

[133] Toshiro Tango. Statistical model of changes in repeated multivariate measurements associated with the development of disease. Computational Statistics & Data Analysis, 3(??): 77–88, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900647.

Wang:1985:ACM

[134] C. Ming Wang. Applications and computing of mosaics. Computational Statistics & Data Analysis, 3 (??):89–97, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900659.

Heilbrun:1985:SSD

[135] Lance K. Heilbrun and Daniel L.

McGee. Sample size determination for the comparison of normal means when one sample size is fixed. Computational Statistics & Data Analysis, 3(??):99–102, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900660.

Matuszewski:1985:BSP

[136] Andrzej Matuszewski and David Sotres.

A basic statistical problem: Confidence interval for the Bernoulli parameter. Computational Statistics & Data Analysis, 3(??):103–114, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900672.

Bailey:1985:NLB

[137] R. A. Bailey. A note on loosely balanced incomplete block designs. Computational Statistics & Data Analysis, 3(??):115–117, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900684. See

[138].

Cromp:1985:RRBa

[138] R. F. Cromp and N. V. Findler.

Rejoinder to R. A. Bailey’s critique “A note on loosely balanced incomplete block designs”. Computational Statistics & Data Analysis, 3(??): 119–120, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900696. See

[138, 137] and response [139].

Bailey:1985:RRF

[139] R. A. Bailey. Response to R. F. Cromp and N. V. Findler. Computational Statistics & Data Analysis, 3(??):121–122, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900702. See

[138] and reply [140].

Cromp:1985:RRBb

[140] R. F. Cromp and N. V. Findler. Reply to R. A. Bailey’s response. Computational Statistics & Data Analysis, 3(??):123–124, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900714. See

[139].

Baker:1985:FCM

[141] Robert J. Baker, Michael R. B. Clarke, and Peter W. Lane. Further comment on M. B. Brown and C. Fuchs, “On maximum likelihood estimation in sparse contingency tables”. Computational Statistics & Data Analysis, 3(??):125–127, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900726. See

[2].

Brown:1985:R

[142] Morton B. Brown and Camil Fuchs.

Rejoinder. Computational Statistics & Data Analysis, 3(??):129– 130, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900738. See

comment [143].

Baker:1985:CRM

[143] R. J. Baker, M. R. B. Clarke, and P. W. Lane. Comment on the rejoinder by M. B. Brown and C. Fuchs. Computational Statistics & Data Analysis, 3(??):131, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590074X. See

[142].

Kullback:1985:MDI

[144] Solomon Kullback and John C. Keegel.

The minimum discrimination information approach to the analysis of contingency tables. Computational Statistics & Data Analysis, 3(??): 133–141, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900751.

Brown:1985:RKK

[145] Morton B. Brown. Rejoinder to Kullback and Keegel note on the minimum discrimination information approach. Computational Statistics & Data Analysis, 3(??):143–144, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900763. See

[142].

Kullback:1985:RMB

[146] S. Kullback and J. C. Keegel. Response to M. B. Brown. Computational Statistics & Data Analysis, 3 (??):145–146, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900775.

Anonymous:1985:Ab

[147] Anonymous. Announcements. Computational Statistics & Data Analysis, 3(??):147–150, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900787.

Gebski:1985:SPS

[148] Val J. Gebski. Some properties of splicing when applied to nonlinear smoothers. Computational Statistics & Data Analysis, 3(??): 151–157, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900799.

Verbeek:1985:SAE

[149] Albert Verbeek and Pieter M. Kroonenberg. A survey of algorithms for exact distributions of test statistics in r × c contingency tables with fixed margins. Computational Statistics & Data Analysis, 3(??):159–185, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900805.

Dodge:1985:MCF

[150] Yadolah Dodge and Kasra Afsarinejad. Minimal 2^{n} connected factorial experiments. Computational Statistics & Data Analysis, 3(??): 187–200, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900817.

Barlow:1985:ABD

[151] William E. Barlow and Polly Feigl. Analyzing binomial data with a nonzero baseline using GLIM. Computational Statistics & Data Analysis, 3 (??):201–204, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900829.

Shea:1985:RSL

[152] Brian L. Shea. The role of a subroutine library for statisticians. Computational Statistics & Data Analysis, 3(??):205–206, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900830.

Anonymous:1985:Ac

[153] Anonymous. Announcements. Computational Statistics & Data Analysis, 3(??):207–218, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900842.

Gupta:1985:ETM

[154] Ramesh Gupta, Ram Tripathi, Joel Michalek, and Thomas White. An exact test for the mean of a normal distribution with a known coefficient of variation. Computational Statistics & Data Analysis, 3(??): 219–226, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900854.

Sadler:1985:RME

[155] W. A. Sadler and M. H. Smith. A reliable method of estimating the variance function in immunoassay. Computational Statistics & Data Analysis, 3(??):227–239, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900866.

Matuszewski:1985:STB

[156] Andrzej Matuszewski and David Sotres. A simple test for the Behrens– Fisher problem. Computational Statistics & Data Analysis, 3(??):241– 249, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900878.

Denteneer:1985:FAI

[157] Dee Denteneer and Albert Verbeek. A fast algorithm for iterative proportional fitting in log-linear models. Computational Statistics & Data Analysis, 3(??):251–264, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590088X.

Neffendorf:1985:SSS

[158] Hugh Neffendorf and David Cable. Software for statistical and survey analysis. Computational Statistics & Data Analysis, 3(??):265–269, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900891.

Anonymous:1985:Ad

[159] Anonymous. Announcements. Computational Statistics & Data Analysis, 3(??):271–282, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947385900908.

Anonymous:1985:AIVb

[160] Anonymous. Author index volume 3 (1985). Computational Statistics & Data Analysis, 3(??):283– 284, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738590091X.

Anonymous:1985:PRRb

[161] Anonymous. Package report/review index volume 3 (1985). Computational Statistics & Data Analysis, 3 (??):285, May 1985. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900921.

Anonymous:1985:EB

[162] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 3(??):IFC, May 1985. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947385900520.

Wixley:1988:UAP

[163] R. A. J. Wixley. The use of asymmetric power transformation in databased quasi-likelihood analysis. Computational Statistics & Data Analysis, 7(1):1–21, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947388900126.

Simeone:1988:FDM

[164] B. Simeone and I. Verdinelli. A feasible directions method for computing Bayes E-optimal block designs. Computational Statistics & Data Analysis, 7(1):23–38, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947388900138.

Draper:1988:BAS

[165] N. R. Draper, M. Evans, and Irwin Guttman. A Bayesian approach to system reliability when two components are dependent. Computational Statistics & Data Analysis, 7 (1):39–49, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738890014X.

Tarter:1988:GED

[166] Michael E. Tarter and William R. Freeman. On graphing estimated distributions using modified scatter diagrams. Computational Statistics & Data Analysis, 7(1):51–67, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900151.

Zelterman:1988:OSG

[167] Daniel Zelterman. Order statistics of the generalized logistic distribution. Computational Statistics & Data Analysis, 7(1):69–77, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900163.

Hsu:1988:SSC

[168] Jason C. Hsu. Sample size computation for designing multiple comparison experiments. Computational Statistics & Data Analysis, 7(1):79– 91, August 1988. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900175.

Blyth:1988:LE

[169] Colin R. Blyth. Letter to the Editor.

Computational Statistics & Data Analysis, 7(1):93–98, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900187.

Anonymous:1988:SCI

[170] Anonymous. The 6th Scandinavian conference on image analysis. Computational Statistics & Data Analysis, 7(1):99–100, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900199.

Anonymous:1988:SAS

[171] Anonymous. SPSS announces SPSS-X for Cray Unicos. Computational Statistics & Data Analysis, 7(1):100– 101, August 1988. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900205.

Anonymous:1988:NEG

[172] Anonymous. New expanded guide of the SAS system for personal computers. Computational Statistics & Data Analysis, 7(1):101–102, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900217.

Anonymous:1988:SCC

[173] Anonymous. SHAZAM: a comprehensive computer program for regression models (version 6). Computational Statistics & Data Analysis, 7 (1):102–104, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900229.

Anonymous:1988:MAV

[174] Anonymous. MicroMath announces version 2.3 of MINSQ. Computational Statistics & Data Analysis, 7 (1):104–105, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900230.

Anonymous:1988:MAG

[175] Anonymous. MicroMath announces GRAPH. Computational Statistics & Data Analysis, 7(1):105– 106, August 1988. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900242.

Bentler:1988:GML

[176] P. M. Bentler, Wai-Yin Poon, and Sik-Yum Lee. Generalized multimode latent variable models: Implementation by standard programs. Computational Statistics & Data Analysis, 7 (2):107–118, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900862.

Kappenman:1988:NBR

[177] Russell F. Kappenman. Nonparametric binary regression without replication. Computational Statistics & Data Analysis, 7(2):119–125, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947388900874.

Coffey:1988:NMV

[178] Mary P. Coffey, Marcia Feingold, and Judith Bromberg. A normed measures of variability among proportions. Computational Statistics & Data Analysis, 7 (2):127–141, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900886.

Hadi:1988:DCI

[179] Ali S. Hadi. Diagnosing collinearityinfluential observations. Computational Statistics & Data Analysis, 7 (2):143–159, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900898.

Little:1988:ACS

[180] Roderick J. A. Little. Approximately calibrated small sample inference about means from bivariate normal data with missing values. Computational Statistics & Data Analysis, 7 (2):161–178, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900904.

Bonett:1988:PEL

[181] Douglas G. Bonett and J. Arthur Woodward. Parameter estimation in linear multinomial models. Computational Statistics & Data Analysis, 7 (2):179–187, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900916.

Bunch:1988:CSE

[182] James R. Bunch. Calculating selected elements of the inverse of symmetric positive definite matrices. Computational Statistics & Data Analysis, 7 (2):189–195, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900928.

Simon:1988:BNA

[183] Stephen D. Simon and James P. Lesage.

Benchmarking numerical accuracy of statistical algorithms. Computational Statistics & Data Analysis, 7(2):197– 209, December 1988. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738890093X.

Anonymous:1988:IPP

[184] Anonymous. IRIS: program package for multivariate analysis. Computational Statistics & Data Analysis, 7 (2):211–212, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900941.

Anonymous:1988:CPE

[185] Anonymous. Call for papers: 1989 European meeting-psychometric society. Computational Statistics & Data Analysis, 7(2):212–213, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/0167947388900953.

Anonymous:1988:ICR

[186] Anonymous. International conference on recent developments in statistical data analysis and inference — in honor of C. R. Rao: 21–24 August 1989, Neuchâtel, Switzerland. Computational Statistics & Data Analysis, 7 (2):213–215, December 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947388900965.

Smith:1989:APS

[187] D. M. Smith and J. M. Bremner. All possible subset regressions using the QR decomposition. Computational Statistics & Data Analysis, 7 (3):217–235, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900236.

Fung:1989:QBE

[188] Wing-Kam Fung. Quasi-Bayesian estimation of Stigler’s data sets. Computational Statistics & Data Analysis, 7 (3):237–243, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900248.

Agresti:1989:MBB

[189] Alan Agresti and Christy Chuang. Model-based Bayesian methods for estimating cell proportions in crossclassification tables having ordered categories. Computational Statistics & Data Analysis, 7(3):245– 258, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738990025X.

Gomes:1989:GGL

[190] M. Ivette Gomes. Generalized Gumbel and likelihood ratio test statistics in the multivariate GEV model. Computational Statistics & Data Analysis, 7 (3):259–267, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900261.

Ritov:1989:MCC

[191] Y. Ritov. Monte Carlo computation of the mean of a function with convex support. Computational Statistics & Data Analysis, 7(3):269– 277, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900273.

Preisler:1989:FDR

[192] Haiganoush K. Preisler. Fitting doseresponse data with non-zero background within generalized linear and generalized additive models. Computational Statistics & Data Analysis, 7 (3):279–290, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900285.

Kalaba:1989:FPT

[193] Robert Kalaba, Nima Rasakhoo, and Leigh Tesfatsion. A FORTRAN program for time-varying linear regression via flexible least squares. Computational Statistics & Data Analysis, 7 (3):291–309, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900297.

Deutler:1989:SRL

[194] T. Deutler. Some remarks on the likelihood ratio test for the mean of a normal distribution with known coefficient of variation. Computational Statistics & Data Analysis, 7(3):311– 314, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900303.

Anonymous:1989:AMS

[195] Anonymous. 13th annual meeting of the society of classification. Computational Statistics & Data Analysis, 7(3):315, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900315.

Anonymous:1989:SCI

[196] Anonymous. Second Conference of the International Federation of Classification Societies: June 27–30, 1989, in Charlottesville, VA, USA. Computational Statistics & Data Analysis, 7(3):316, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900327.

Anonymous:1989:FAC

[197] Anonymous. First announcement and call for papers — joint meeting: GLIM 89 and 4^{th} international workshop on statistical modelling: Trento, Italy, July 17–21, 1989. Computational Statistics & Data Analysis, 7 (3):317–318, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900339.

Anonymous:1989:SISa

[198] Anonymous. SPSS introduces SPSS/PC + v3.0. Computational Statistics & Data Analysis, 7(3):318– 319, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900340.

Anonymous:1989:TSP

[199] Anonymous. Targets scientists for PC data analysis. Computational Statistics & Data Analysis, 7(3):319– 320, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900352.

Anonymous:1989:SV

[200] Anonymous. SYSTAT version 4.0. Computational Statistics & Data Analysis, 7 (3):320–321, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900364.

Anonymous:1989:S

[201] Anonymous. Sygraph. Computational Statistics & Data Analysis, 7 (3):321–322, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900376.

Anonymous:1989:SVM

[202] Anonymous. SYSTAT version 3.2 for Macintosh. Computational Statistics & Data Analysis, 7(3):322– 323, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900388.

Anonymous:1989:BSSa

[203] Anonymous. BMDP statistical software — latest release. Computational Statistics & Data Analysis, 7 (3):323–325, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794738990039X.

Anonymous:1989:RM

[204] Anonymous. La revue de modulad. Computational Statistics & Data Analysis, 7(3):325–326, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/0167947389900406.

Anonymous:1989:AIVa

[205] Anonymous. Author index volume 7 (1989). Computational Statistics & Data Analysis, 7(3):327– 328, February 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900418.

Anonymous:1989:PRRa

[206] Anonymous. Package report/review index volume 7 (1989). Computational Statistics & Data Analysis, 7 (3):329, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738990042X.

Anonymous:1989:EB

[207] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 8(1):ii, May 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900583.

Anonymous:1989:P

[208] Anonymous.  Preface.  Computa-

tional Statistics & Data Analysis, 8 (1):v, May 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900595.

Pecar:1989:COC

[209] Ján Pecár. Combinations of optimality criteria and optimum designs. Computational Statistics & Data Analysis, 8(1):1–11, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900601.

Kitsos:1989:FSP

[210] C. P. Kitsos. Fully-sequential procedures in nonlinear design problems. Computational Statistics & Data Analysis, 8(1):13–19, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900613.

Giovagnoli:1989:EDM

[211] Alessandra Giovagnoli and Paola Sebastiani. Experimental designs for mean and variance estimation in variance components models. Computational Statistics & Data Analysis, 8 (1):21–28, May 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900625. See

comments [252].

Gupta:1989:ENE

[212] Sudhir Gupta and Rahul Mukerjee. Efficient non-equireplicate designs obtained by merging of treatments. Computational Statistics & Data Analysis, 8(1):29–37, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900637.

Kaishev:1989:OED

[213] V. K. Kaishev. Optimal experimental designs for the B-spline regression. Computational Statistics & Data Analysis, 8(1):39–47, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900649.

Li:1989:AOR

[214] Kim-Hung Li and N. N. Chan. An algorithm for A-optimal regression designs. Computational Statistics & Data Analysis, 8(1):49–57, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900650.

Bhaumik:1989:MON

[215] Dulal Kumar Bhaumik and Thomas Mathew. On the minimax optimality of non-randomized block designs. Computational Statistics & Data Analysis, 8(1):59–66, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900662.

Tricot:1989:RTA

[216] M. L. Tricot. Randomization tests in ANOVA: a practical approach. Computational Statistics & Data Analysis, 8(1):67–74, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900674.

Ermakov:1989:RIT

[217] Sergei M. Ermakov. Random interpolation in the theory of experimental design. Computational Statistics & Data Analysis, 8(1): 75–80, May 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900686.

Kettunen:1989:ASD

[218] Juhani Kettunen, Hannu Sirviö, and Olli Varis. An analysis of a sampling design — a case study of lake eutrophication. Computational Statistics & Data Analysis, 8(1):81–91, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900698.

vanBerkum:1989:RNP

[219] E. E. M. van Berkum. Reduction of the number of pairs in paired comparison designs and exact designs for quadratic models. Computational Statistics & Data Analysis, 8(1):93–107, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900704.

Wang:1989:TDE

[220] P. C. Wang. Test for dispersion effects from orthogonal arrays. Computational Statistics & Data Analysis, 8(1):109–117, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900716.

Wierich:1989:REB

[221] W. Wierich. On the relative efficiency of the best second order proportional ANOVA-designs. Computational Statistics & Data Analysis, 8(1):119–134, May 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900728.

Chakravarti:1989:BIM

[222] Nilotpal Chakravarti. Bounded isotonic median regression. Computational Statistics & Data Analysis, 8 (2):135–142, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900017.

Frydenberg:1989:MIP

[223] Morten Frydenberg and Davis Edwards. A modified iterative proportional scaling algorithm for estimation in regular exponential families. Computational Statistics & Data Analysis, 8(2):143–153, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900029.

Marron:1989:CDB

[224] J. S. Marron. Comments on a data based bandwidth selector. Computational Statistics & Data Analysis, 8(2):155–170, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900030.

Turek:1989:ETG

[225] Richard J. Turek and Ronald C.

Suich. An exact test on the Goodman– Kruskal λ for prediction on a dichotomy: power considerations. Computational Statistics & Data Analysis, 8(2):171–178, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900042.

Gray:1989:FMI

[226] J. Brian Gray. The four-measure influence plot. Computational Statistics & Data Analysis, 8(2):179– 188, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900054.

Merckens:1989:FMS

[227] Arjen Merckens and Tom Wansbeek. Formula manipulation in statistics on the computer: Evaluating the expectation of higher-degree functions of normally distributed matrices. Computational Statistics & Data Analysis, 8(2):189–200, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900066.

Schwertman:1989:SAS

[228] Neil C. Schwertman and Margaret A. Owens. Simple approximation of sample size for the bivariate normal. Computational Statistics & Data Analysis, 8(2):201–207, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900078.

Anonymous:1989:LE

[229] Anonymous. Letter to the editor. Computational Statistics & Data Analysis, 8(2):209, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738990008X.

Anonymous:1989:E

[230] Anonymous. Erratum. Computational Statistics & Data Analysis, 8 (2):210, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900091.

Anonymous:1989:SAI

[231] Anonymous. SPSS announces interfaces to Informix and Sybase. Computational Statistics & Data Analysis, 8(2):211, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900108.

Anonymous:1989:SII

[232] Anonymous. SPSS Inc. introduces SPSS Graphics with databases interfaces for Informix and Sybase. Computational Statistics & Data Analysis, 8(2):212, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016794738990011X.

Anonymous:1989:SIV

[233] Anonymous. SPSS increases value of SPSS/PC + by bundling its graphin-the-box option. Computational Statistics & Data Analysis, 8(2): 212–213, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900121.

Anonymous:1989:GRD

[234] Anonymous. Genstat 5 release 1.3 for Dec Vax VMS and MicroVax VMS. Computational Statistics & Data Analysis, 8(2):213–214, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900133.

Anonymous:1989:BSSb

[235] Anonymous. BMDP statistical software introduces Solo: the stand-alone solution for non-statisticians. Computational Statistics & Data Analysis, 8(2):214–216, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900145.

Anonymous:1989:BSSc

[236] Anonymous. BMDP statistical software ships new PC release. Computational Statistics & Data Analysis, 8(2):216–217, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900157.

Anonymous:1989:BSSd

[237] Anonymous. BMDP statistical software announces latest release with two new programs. Computational Statistics & Data Analysis, 8(2): 218–219, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900169.

Anonymous:1989:SISb

[238] Anonymous. SYSTAT, Inc. software for data general and NCR/TOWER. Computational Statistics & Data Analysis, 8(2):219–220, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900170.

Anonymous:1989:SIA

[239] Anonymous. SYSTAT, Inc. announces version 1.1 of SYGRAPH at NCGA booth 4457. Computational Statistics & Data Analysis, 8 (2):220–221, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900182.

Anonymous:1989:SWO

[240] Anonymous. Stat wars: an overview of statistical software and SYSTAT, Inc. Computational Statistics & Data Analysis, 8(2):221–222, July 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900194.

Anonymous:1989:APS

[241] Anonymous. Announcing powerful software for solving ordinary differential equations. Computational Statistics & Data Analysis, 8(2): 223–224, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900200.

Anonymous:1989:CTT

[242] Anonymous. Connectivity tool transfers data among database and statistical products. Computational Statistics & Data Analysis, 8(2): 224, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900212.

Anonymous:1989:AUD

[243] Anonymous. Analele Universităţii Din Bucureşti. Computational Statistics & Data Analysis, 8(2):224– 225, July 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900224.

DeVeaux:1989:MLR

[244] Richard D. De Veaux. Mixtures of linear regressions. Computational Statistics & Data Analysis, 8(3):227– 245, November 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900431.

Devroye:1989:RVG

[245] Luc Devroye. Random variate generators for the Poisson — Poisson and related distributions. Computational Statistics & Data Analysis, 8 (3):247–278, November 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900443.

Gleason:1989:ESD

[246] John R. Gleason. Efficient sorting during repetitive statistical computations: algorithms and an application. Computational Statistics & Data Analysis, 8 (3):279–297, November 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900455.

Malvestuto:1989:CME

[247] F. M. Malvestuto. Computing the maximum-entropy extension of given discrete probability distributions. Computational Statistics & Data Analysis, 8 (3):299–311, November 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900467.

Shih:1989:HPE

[248] Wei Shih and James A. Sullivan. A heuristic parameter estimation procedure for a binary dependent variable regression model. Computational Statistics & Data Analysis, 8(3):313– 324, November 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167947389900479.

Simon:1989:AAA

[249] Stephen D. Simon and James P. Lesage.

Assessing the accuracy of ANOVA calculations in statistical software. Computational Statistics & Data Analysis, 8 (3):325–332, November 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900480.

Anonymous:1989:AIVb

[250] Anonymous. Author index volume 8 (1989). Computational Statistics & Data Analysis, 8(3):333– 334, November 1989. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900492.

Anonymous:1989:PRRb

[251] Anonymous. Package report/review index volume 8 (1989). Computational Statistics & Data Analysis, 8 (3):335, November 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167947389900509.

Anonymous:1990:EDM

[252] Anonymous. “Experimental designs for mean and variance estimation in variance components models”: by Alessandra Giovagnoli and Paola Sebastiani, Computational Statistics & Data Analysis 8, 21–28. Computational Statistics & Data Analysis, 10 (1):93, August 1990. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016794739090106R. See

[211].

Francis:1981:SSC

[253] Ivor Francis. Statistical software: a comparative review. North-Holland Publishing Co., Amsterdam, The Netherlands, 1981. ISBN 0-444-00658-3. xx + 542 pp. LCCN QA276.4 .F7 1981.

David:1982:CPS

[254] Hugh David. The computer package STATCAT: source programs and user manual. North-Holland Publishing Co., Amsterdam, The Netherlands, 1982. ISBN 0-444-86453-9. ix + 780 pp. LCCN QA276.4 .D38 1982.

Lewi:1982:MDA

[255] Paul J. Lewi. Multivariate data analysis in industrial practice. Chemometrics research studies series. Research Studies Press, Chichester, West Sussex, UK, 1982. ISBN 0-471-10466-3. xi + 244 pp. LCCN QA278 .L48 1982. US$41.30.